# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v1

@@ 1.3 Definitions
The metrics under consideration in this article are the Mean-Median Difference and the Partisan Bias. These metrics are well-known in the community (see, for example, [15]), but for completeness we give their definitions here. We shall also show how they can be visualized on the seats-votes curve.
@@
In the next couple of sections, (Sections 1.4 and 1.5) we assume that the statewide vote share $V$ is simply the average of the vote shares in each district: $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. In particular, note that the creation of the seats-votes curve assumes that the statewide vote share $V$ satisfies $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This is true when turnout in each district is the same (though is not true in general). We drop this assumption in Section 1.6.
@@ 1.4 The Seats-Votes Curve
Both the Mean-Median Difference and the Partisan Bias can be visualized on the seats-votes curve corresponding to the election with data $V_{1},V_{2},\dots,V_{n}$. The seats-votes curve is a curve drawn in the vote-share/seat-share plane under the assumption of uniform partisan swing. More specifically, the data for a single election corresponds to a vote share $V$ (for now assumed to be the mean of the $V_{i}$s), and a seat share $S$. One assumes that the vote shares $V_{i}$ swing towards party $A$ in a uniform manner; they all increase by the same vote share or decrease by the same vote share. This results in a new vote share $V^{\prime}$, and (depending on whether that vote share swing was sufficient to flip a district) a potentially new seat share $S^{\prime}$. All such new $(V^{\prime},S^{\prime})$ are drawn in the plane, which results in a curve which is a step function. Two examples can be seen in Figure 1 below. One curve corresponds to the seats-votes curve for a map whose vote shares are 0.2, 0.3, 0.55, 0.6, and 0.65; the other corresponds to real data from the Massachusetts 2011 congressional map.
@@
Please note that the correspondence between the $MM$, $PB$, and the seats-votes curve described above is only true when one can reach $S=0$ and $S=1$ under uniform partisan swing. That is, this is assuming that in order to reach $S=1$ (or $S=0$) under uniform partisan swing, we don’t push a district’s vote share above 1 (or below 0) [^3] If $V_{1}$ is party $A$’s minimum vote share in a district and $V_{n}$ is party $A$’s maximum vote share in a district, then we can mathematically state this as follows: $0.5-V_{1}<1-V_{n}$ and $V_{n}-0.5<V_{1}$.
Of course, uniform partisan swing is by no means guaranteed or even likely to happen in practice (see [37] for other models of swing). Nevertheless, metrics derived from the seats-votes curve, referred to as partisan symmetry metrics, were the first election outcome metrics introduced to detect partisan gerrymandering, and have had significant staying power (as mentioned in Section 1). Indeed, the authors of [10] pointed out amicus briefs that have used partisan symmetry metrics, including LULAC v. Perry (2006) [24], Whitford v. Gill (2018) [14], and Rucho v. Common Cause (2019) [17]. Partisan symmetry metrics have also been used in various expert testimony reports, including Session v. Perry (2003) [2], Perez v Perry (2011) [27], Favors v. Cuomo (2012) [3], and an appeal of Clarke v. Wisconsin Elections Commission (2023) [9]. We hope that with articles such as [10], [30], and this article, experts will use symmetry metrics such as the Mean-Median Difference and Partisan Bias with significantly more caution.
@@ 1.5 Relationship between Mean-Median Difference and Partisan Bias
The Mean-Median Difference and Partisan Bias metrics cannot disagree in sign: if one is positive, the other is non-negative, and if one is negative the other is not positive. This follows directly from the fact that the seats-votes curve is a non-decreasing curve going through points (0,0) and (1,1). Note that whenever $MM=0$, the median of the $V_{i}$s is equal to the mean. This implies that the number of $V_{i}$ above the mean must be equal to the number of $V_{i}$s below; in other words, it implies that $PB=0$.
@@
The supremal difference between $PB$ and $MM$ is 0.5. This can be argued algebraically, but it can be seen more easily by considering how the $MM$ and $PB$ correspond to points on the seats-votes curve. For example, if $MM<0$, then the supremal value of $MM-PB$ occurs when $PB$ is near -0.5 and $MM$ is near 0. An example of such a seats-votes curve can be seen in Figure 2.
@@ 1.6 Unequal Turnout
As we shall see in Section 2, if we consider election data corresponding to uneven turnout between districts, we can get a wider array of $(V,S)$ pairs giving $MM=0$ and $PB=0$.
2 Theoretical Results
We first note that, if an election has vote-share $V$ and seat-share $S$, then we must have $0\leq V\leq 1,0\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$ (see, for example, [33]).
Please note that many of our constructed election data cannot be visualized on the seats-votes curve under the assumption of uniform partisan swing. (Recall from Section 1.4 that this may be the case if, under uniform partisan swing, we must have some districts with vote share less than 0 in order to achieve $S=0$ and/or districts with vote share more than 1 in order to achieve $S=1$). Also please note that our constructed data does not limit the number of districts. The main theoretical results of this article are below.
@@ 2 Theoretical Results
Note that if party $A$’s seat share is $S$, then $1-S$ is party $B$’s seat share, and that party $A$’s PB value (and MM value) is the negation of party $B$’s PB value (MM value). Thus, Theorem 1 gives all possible range of values for PB for any possible $(V,S)$. These minimum and maximum values can be visualized in Figure 3. Theorem 1 is proved in Appendix A.
@@
In Figure 3, we can visually see a discontinuity at $V=\frac{1}{2}$. We can understand this intuitively as well, by using the following general argument: fix $V\not=\frac{1}{2}$ and $S\geq\frac{1}{2}$. Then we can construct election data with the correct number of winning and losing districts such that the number of districts whose vote share is equal to $V$ is maximized. If we barely raise the vote shares of nearly all of the districts whose vote share is $V$ (and lower a single seat share), this maximizes PB, and keeps the correct vote share and seat share. If we barely lower the vote shares of nearly all of the districts whose vote share is $V$ (and raise a single seat share), this minimizes PB and keeps the correct vote share and seat share.
However, this argument breaks down when $V=\frac{1}{2}$ because barely raising or lowering the vote share can change it from winning to losing or vice versa. This creates the discontinuity at that line.
We also have the corresponding Theorem for the Mean-Median Difference:
@@
From Figures 3 and 4, we can see that the range of values that MM and PB can take respond differently to changing $V$ and $S$. Consider, for example, the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{2}\leq V\leq 1$. For a fixed $S$ in that region, the maximum possible value for $PB$ doesn’t depend on $V$ at all, as it is $S-\frac{1}{2}$. Additionally, note that for $V$ just barely less than $\frac{1}{2}$, PB can achieve nearly the maximal value of $\frac{1}{2}$, regardless of $S$. (And symmetrically, for $V$ just barely above $\frac{1}{2}$, PB can achieve nearly the minimal value of $-\frac{1}{2}$).
For the Mean-Median Difference, we can consider the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{4}\leq V\leq\frac{3}{4}$. In this region, for a fixed $V$, regardless of the value of $S$, the Mean-Median Difference has the same possible minimum value ($\frac{1}{2}-V$), while the largest maximum values occur near the line $S=2V-\frac{1}{2}$, which happens to be the line where the Efficiency Gap is always 0. Indeed, the Efficiency Gap can only take on a single value when turnout in each district is equal, which can be seen in Figure 5.
Figure 5: Only possible value of Efficiency Gap (using constructed election data) for each vote-share, seat-share pair $(V,S)$, when turnout in each district is equal.
While these observations may be mathematically interesting, some of them seem potentially problematic in a metric intended to detect partisan gerrymandering. For example, when considering the Partisan Bias: when $S$ with $\frac{1}{2}\leq V\leq 1$ is fixed, but $V$ increases, why should PB be able to take on values indicating the same amount of benefit to the party in question? And why should nearly the maximum benefit be able to be indicated for $V$ just under $\frac{1}{2}$ and $S$ any value? For the Mean-Median Difference: for fixed $V$, why would the largest possible values be achievable on the line $S=2V-\frac{1}{2}$?
These questions are unanswerable, as these mathematical effects are surely not what was considered in the design of the Mean-Median Difference and Partisan Bias. Nevertheless, we pose them simply to point out the likely unintended consequences of the definitions of MM and PB.

### v3

@@ 3 Definitions and other Established Facts
The Mean-Median Difference and Partisan Bias metrics are well-known in the community, but for completeness we give their definitions here. We shall also share other established facts about these metrics, to help the unfamiliar reader to form an intuition and deeper understanding of these metrics.
@@
In the next couple of sections, (Sections 3.1 and 3.2) we assume that the statewide vote share $V$ is simply the average of the vote shares in each district: $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. In particular, note that the creation of the seats-votes curve assumes that the statewide vote share $V$ satisfies $V=\frac{1}{n}\sum_{i=1}^{n}V_{i}$. This is true when turnout in each district is the same (though is not true in general). We drop this assumption in Section 3.3.
@@ 3.1 The Seats-Votes Curve
Both the Mean-Median Difference and the Partisan Bias can be interpreted in terms of the seats-votes curve when drawn under the assumption of uniform partisan swing. It is worth noting that there are many different ways to draw a seats-votes curve; see Section 1.1.1 of [15] for a nice summary of various methods. The construction of the seats-votes curve using uniform partisan swing starts with estimates for the party $A$ vote share $V_{i}$ in each district $i=1,2,\dots,n$. This data corresponds to a statewide vote share $V$ (for now assumed to be the mean of the $V_{i}$s), and seat share $S$. One assumes that if the statewide vote share swings towards party $A$, the vote shares $V_{i}$ swing towards party $A$ in a uniform manner; they all increase by the same vote share or decrease by the same vote share. This results in a new vote share $V^{\prime}$, and (depending on whether that vote share swing was sufficient to flip a district) a potentially new seat share $S^{\prime}$. All such new $(V^{\prime},S^{\prime})$ are drawn in the plane, which results in a curve which is a step function. Two examples can be seen in Figure 1 below. One curve corresponds to the seats-votes curve for a map whose vote shares are 0.2, 0.3, 0.55, 0.6, and 0.65; the other corresponds to real data from the Massachusetts 2011 congressional map.
@@
The correspondence between the $MM$, $PB$, and the seats-votes curve described above is only true when the curve is drawn under the assumption of uniform partisan swing. Of course, uniform partisan swing is by no means guaranteed or even likely to happen in practice (see [52] for other models of swing). Grofman and King [29] do not even define the Partisan Bias as we do in Definition 2, and instead define it using the intersection of the seats-votes curve with the line $V=0.5$, where the seats-votes curve is constructed not by assuming uniform partisan swing, but by using statistical methods focused on the rank order [^5] Rank is in terms of partisan support for a particular political party. of districts. Grofman and King acknowledge that Partisan Bias is frequently defined as in our Definition 2, but call the definition of the Partisan Bias which uses the aforementioned statistical technique of defining the seats-votes curve the “state of the art” version of the metric [29]. However, as DeFord et al point out, complex definitions (such as Grofman and King’s “state of the art” version of Partisan Bias) must be “translated to the real-world ‘political thicket’ of redistricting” [16]. Wang agrees, by stating that a manageable standard to detect partisan gerrymandering “should be able to be clearly stated without case-specific or mathematics-intensive assumptions” [49].
Indeed, our Definitions 1 and 2 are precisely the implementations of Mean-Median Difference and Partisan Bias that are currently being used in the “real-world political thicket of redistricting.” Dave’s Redistricting App and Maptitude both calculate the Mean-Median Difference and Partisan Bias identically to Definitions 1 and 2, subject to choice of underlying vote data. Similarly, the Princeton Gerrymandering Project’s Report Cards [50] use the same type of calculation and the Python library GerryChain [40] has built-in methods for calculating the Mean-Median Difference and Partisan Bias, again using formulations identical to our Definitions 1 and 2. PlanScore [26] also reports values of the Mean-Median Difference and Partisan Bias, contributing to the ubiquity of these metrics [^6] The definitions, vote data, and metric scores on PlanScore’s site are less clear about the exact aproach used by Planscore to calculate the MM and PB. These calculations are also ubiquitous in academic research and expert testimony. We believe it not only useful, but imperative to study the version of the Partisan Bias (and Mean-Median Difference) that is used in software, websites analyzing maps, research, and expert testimony, which is why we define those metrics as they are being used there.
@@ 3.2 Relationship between Mean-Median Difference and Partisan Bias
The Mean-Median Difference and Partisan Bias metrics cannot disagree in sign: if one is positive, the other is non-negative, and if one is negative the other is not positive. This follows directly from the ways the MM and PB can be calculated on a seats-votes curve drawn using uniform partisan swing (as described in Section 3.1), and the fact that the seats-votes curve is a non-decreasing curve going through points (0,0) and (1,1). Note that whenever $MM=0$, the median of the $V_{i}$s is equal to the mean. This implies that the number of $V_{i}$ above the mean must be equal to the number of $V_{i}$s below; in other words, it implies that $PB=0$.
@@
The supremal difference between $PB$ and $MM$ is 0.5. This can be argued algebraically, but it can also be seen by considering how the $MM$ and $PB$ correspond to points on the seats-votes curve. For example, if $MM<0$, then the supremal value of $MM-PB$ occurs when $PB$ is near -0.5 and $MM$ is near 0. An example of such a seats-votes curve can be seen in Figure 2.
@@ 3.3 Unequal Turnout
As we shall see in Section 4, if we consider election data corresponding to uneven turnout between districts, we can get a wider array of $(V,S)$ pairs giving $MM=0$ and $PB=0$. The impact of turnout differentials on measures of partisan fairness can have significant impacts on their interpretation and robustness, as discussed for the Efficiency Gap in [48].
4 Bounds
Here, we give our mathematical results showing which values the Mean-Median Difference and Partisan Bias can take on for a fixed vote-share, seat-share pair $(V,S)$. As we shall see, a fixed $(V,S)$ pair can correspond to an array of different data, which give vastly different values for a single metric (either the Mean-Median Difference or Partisan Bias). Thus, our results will show that a map could have a fixed vote share, a large seat share, and the metric could indicate that the party with large seat share is gerrymandered against. And the same map could have the same fixed vote share, low seat share, and the metric could indicate that the map is gerrymandered in favor of the party with low seat share. In other words, these results will indicate, at least on constructed data, that these metrics do not indicate when a party wins an extreme number of districts, as suggested by the use cases discussed in Section 2.2.
Of course, real maps and real data may be different from the space of all possible constructed data. This is why we have the empirical arguments in forthcoming Section 5. But for now, we begin the setup of our theory argument. The rigorous mathematical statements of our results are in Sections 4.1 and 4.2. We present a discussion of the implications of these results in Section 4.3.
We first note that, if an election has vote-share $V$ and seat-share $S$, and if we assume turnout in all districts is equal (as we do in Sections 4.1, 4.2, and 4.3) then we must have $0\leq V\leq 1,0\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$ (see, for example, [47]).
Please note that Theorems 1 and 3 have no restrictions on the number of districts. The corresponding Theorems for a fixed number of districts are Theorems 2 and 4.
4.1 Bounds: Partisan Bias
@@
Note that if party $A$’s seat share is $S$, then $1-S$ is party $B$’s seat share, and that party $A$’s PB value is the negation of party $B$’s PB value. Thus, Theorem 1 gives all possible range of values for PB for any possible $(V,S)$. These minimum and maximum values can be visualized in Figure 3. Theorem 1 is proved in Appendix A.
@@
In Figure 3, we can visually see a discontinuity at $V=\frac{1}{2}$. We can understand this intuitively, by using the following general argument: fix $V\not=\frac{1}{2}$ and $S\geq\frac{1}{2}$. Then we can construct election data with the correct number of winning and losing districts such that the number of districts whose vote share is equal to $V$ is maximized. If we barely raise the vote shares of nearly all of the districts whose vote share is $V$ (and lower a single seat share), this maximizes PB, and keeps the correct vote share and seat share. If we barely lower the vote shares of nearly all of the districts whose vote share is $V$ (and raise a single seat share), this minimizes PB and keeps the correct vote share and seat share.
However, this argument breaks down when $V=\frac{1}{2}$ because barely raising or lowering the vote share can change it from winning to losing or vice versa. This results in the discontinuity at that line.
Theorem 1 gives bounds on the value of the Partisan Bias when the number of districts is unlimited. However, using the same arguments as in the proofs in Appendix A, we have the following bounds for a fixed number of districts:
Theorem 2.
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n$ is the number of districts and that $\ell$ is the number of districts lost by the party in consideration, so that $S=\frac{n-\ell}{n}$. Then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
[…]
Again, by the symmetry of the parties in their corresponding PB values, Theorem 2 gives all possible range of values for PB for any possible $(V,S)$. Of course, if we take the limit as $n\to\infty$ in Theorem 2, we obtain the result in Theorem 1.
4.2 Bounds: Mean-Median Difference
We also have corresponding Theorems for the Mean-Median Difference:
@@ 4.2 Bounds: Mean-Median Difference
Again, note that if party $A$’s seat share is $S$, then $1-S$ is party $B$’s seat share, and that party $A$’s MM value is the negation of party $B$’s MM value so that Theorem 3 gives all possible range of values for MM for any possible $(V,S)$.
@@
Theorem 3 gives bounds on the value of the Mean-Median Difference when the number of districts is unlimited. However, using the same arguments as in the proofs in Appendix A, we have the following bounds for a fixed number of districts:
Theorem 4.
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n$ is the number of districts and that $\ell$ is the number of districts lost by the party in consideration, so that $S=\frac{n-\ell}{n}$.
If $S=\frac{1}{2}$, then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Mean-Median Difference $MM$ for any $MM$ in the following ranges:
[…]
If $S>\frac{1}{2}$ and $n$ is even, then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Mean-Median Difference $MM$ for any $MM$ in the following ranges:
[…]
And if $S>\frac{1}{2}$ and $n$ is odd, then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Mean-Median Difference $MM$ for any $MM$ in the following ranges:
[…]
Again, by the symmetry of the parties in their corresponding MM values, Theorem 4 gives all possible range of values for MM for any possible $(V,S)$.
4.3 Bounds: Discussion
From Figures 3 and 4, we can see that the values that MM and PB can take respond differently to changing $V$ and $S$. In this section, we illustrate how the values that Partisan Bias and Mean-Median Difference are able to take on do not generally become more extreme as the number of seats won becomes more extreme (in Sections 4.3.1 and 4.3.3). In sections 4.3.4 and 4.3.2, we also give some specific election examples (and their corresponding $V$, $S$, and metric values) that run counter to expectations of MM and PB tracking extreme seats outcomes, based on the use cases outlined in Section 2.2.
4.3.1 Partisan Bias: Decoupling of Extreme Values with Extreme Number of Districts Won
For a fixed vote share $V$, the Partisan Bias does not in general allow for more extreme values when the seat share $S$ becomes small. Consider a vote share $V>0.5$ for party $A$, and the potential values that the Partisan Bias can take on with different seat shares $S$. From Theorem 1 (and Figure 3), we can see that as $S$ decreases, the minimum values that the Partisan Bias function can take on increase. In other words, for a fixed $V>0.5$, as the number of seats that party $A$ wins decreases the Partisan Bias function is less able to indicate that party $A$ is winning extremely few seats. In particular, if $V>0.75$, once the number of seats won decreases far enough, the Partisan Bias can only take on positive values. This can be seen in the “red crescent” in Figure 3. That is, once the number of seats won by party $A$ is low enough, the Partisan Bias can only indicate a gerrymander in favor of party $A$. The Partisan Bias is not able to indicate that party $A$ is winning extremely few seats [^7] By symmetry, one can similarly show that for $V<0.5$, the Partisan Bias is not able to indicate that party $A$ is winning an extreme number of seats..
Similarly, for a fixed seat share $S$, the Partisan Bias does not allow for more extreme values indicating a less favorable position for party $A$ as party $A$’s vote share $V$ increases, nor does it allow for more extreme values indicating a more favorable position for party $A$ as $V$ decreases. For example, consider a fixed $S$ for party $A$. From Theorem 1 (and Figure 3), we can see that as $V$ increases beyond 0.5, the minimum values that the Partisan Bias can take on increase. That is, for a fixed seat share, if the vote share increases, the Partisan Bias is less able to indicate a gerrymander against party $A$. Indeed, if $S>0.5$ and the vote share increases large enough, the Partisan Bias is only able to take on positive values (the “red crescent” in Figure 3), indicating a gerrymander favoring party $A$. Similarly, for a fixed $S$ for party $A$, as $V$ decreases below 0.5, the maximum values that the Partisan Bias can take on decrease. That is, for a fixed seat share, if the vote share decreases, the Partisan Bias is less able to indicate a gerrymander favoring party $A$.
4.3.2 Partisan Bias: Elections with Values Counter to Public Expectations
In Section 2.2, we mentioned many instances in which the Partisan Bias has been used (or its usage has been promoted) in an effort to avoid maps that allow for one party to win an extreme number of districts. In this section, we give sample elections along with their Partisan Bias calculations. We believe these give convincing evidence that the Partisan Bias does not act in the ways suggested by many of the testimonies, apps, reports, and policies mentioned in Section 2.2.
From Figure 3, we can see that the same coordinate position $(V,S)$ can have extremely different hues in the two images. This corresponds to the fact that, for fixed $V$ and $S$ values, the Partisan Bias can vary wildly. For example, consider the election data in Table 1
District 1 2 3 4 5 6 7 8 9 10
Vote share Election 1
37% 61% 61% 61% 61% 61% 61% 61% 61% 75%
Election 2
49% 59% 59% 59% 59% 59% 59% 59% 59% 79%
Table 1: Two different elections with $S=90\%$ and $V=60\%$. Election 1 has $\text{PB}=0.4$, while Election 2 has $\text{PB}=-0.4$.
Although both Elections in Table 1 have a vote share of 60% and seat share of 90%, the Partisan Bias metric considers Election 1 to be an extreme gerrymander in favor of the party in question, while Election 2 is considered to be an extreme gerrymander against the party in question. The different values here indicate that the Partisan Bias believes that 9 out of 10 seats is extremely large in the first election, but not nearly large enough in the second.
@@ 4.3.2 Partisan Bias: Elections with Values Counter to Public Expectations
Additionally, note in Figure 3 that for $V$ just barely less than $\frac{1}{2}$, PB can achieve nearly the maximal value of $\frac{1}{2}$, regardless of $S$. (And symmetrically, for $V$ just barely above $\frac{1}{2}$, PB can achieve nearly the minimal value of $-\frac{1}{2}$). For example, consider the elections in Table 2.
District 1 2 3 4 5 6 7 8 9 10
Vote share Election 1
35% 49% 49% 49% 49% 49% 49% 49% 49% 53%
Election 2
19% 49% 49% 49% 49% 53% 53% 53% 53% 53%
Election 3
11% 51% 51% 51% 51% 53% 53% 53% 53% 53%
Table 2: Three different elections with $V=48\%$ and $\text{PB}=0.4$. Election 1 has $S=10\%$, Election 2 has $S=50\%$, and Election 3 has $S=90\%$.
In each of the elections in Table 2, the vote share is 48% and the Partisan Bias is 0.4, indicating that each is a very strong gerrymander in favor of the party in question, even though the seat shares vary wildly (from 10% to 50% to 90%). The Partisan Bias is indicating that, for each of the elections in Table 2 (which have the party in question winning 1, 5, and 9 seats out of 10 respectively), an extremely large number of districts is won. Indeed, since $\text{PB}=0.4$ for each of those elections, the Partisan Bias is indicating that 1, 5, and 9 seats won out of 10 are equally extremely large.
The last observation we make in this section is that for a fixed $S$ in the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{2}\leq V\leq 1$, the maximum possible value for $PB$ doesn’t depend on $V$ at all, as it is $S-\frac{1}{2}$. For example, consider the election data in Table 3:
District 1 2 3 4 5 6 7 8 9 10
Vote share Election 1
49% 49% 49% 49% 51% 51% 51% 51% 51% 51%
Election 2
49% 49% 49% 49% 95% 95% 95% 95% 95% 95%
Table 3: Two different elections with $S=60\%$ and $\text{PB}=0.1$. Election 1 has vote share 50.2%, while Election 2 has vote share 76.6%.
We see that, according to the Partisan Bias, these elections are both biased towards the party in question. Moreover, since PB is equal for these two elections, they are equally biased towards the party receiving 6 seats in both elections, even though in Election 1 the party receives less than proportionally many seats and in Election 2 the party receives more than proportionally many seats.
4.3.3 Mean-Median Difference: Decoupling of Extreme Values with Extreme Number of Districts Won
For a fixed vote share $V$ with $0.750.5<V<$ 0.75, the Mean-Median Difference does not in general allow for more extreme values when the seat share $S$ becomes small. Fix a vote share $0.750.5<V<$ 0.75 for party $A$. From Theorem 3 (and Figure 4), we can see that, as party $A$’s number of seats won decreases, the minimum value that the Mean-Median Difference can take on does not change until $S$ crosses from just above 0.5 to just below. Once the seat share $S$ crosses below 0.5, the minimum value that the Mean-Median Difference can take on increases. That is, for a fixed $V$ with $0.750.5<V<$ 0.75 for party $A$ and $S<0.5$, if party $A$ wins fewer seats, the Mean-Median Difference is less able to detect that party $A$ is winning extremely few districts [^8] By symmetry, if $0.50.25<V<$ 0.5, if party $A$ wins more seats, the Mean-Median Difference is less able to detect that party $A$ is winning an extreme number of districts..
Additionally, for a fixed vote share $V$ with $0.750.5<V<$ 0.75, the Mean-Median Difference does not in general allow for more extreme values when the seat share $S$ becomes large. If we fix $V$ with $0.750.5<V<$ 0.75 and now consider the maximum values that the Mean-Median difference can take on, from Theorem 3 and Figure 4 we can see that as party $A$ wins more seats, this maximum value increases to a point and then decreases. In other words, as Party $A$ wins more seats, the Mean-Median Difference can detect that party $A$ is winning a more extreme number of seats . . . until party $A$ wins even more seats, at which point the Mean-Median Difference is less able to detect that party $A$ is winning an extreme number of seats [^9] By symmetry, if $0.50.25<V<$ 0.5, as Party $A$ wins fewer seats, the Mean-Median Difference can detect that party $A$ is winning extremely few seats . . . until party $A$ wins even fewer seats, at which point the Mean-Median Difference is less able to detect that party $A$ is winning extremely few seats..
Finally, for a fixed vote share $V$ with $V>0.75$, the Mean-Median Difference does not in general allow for more extreme values when the seat share $S$ becomes small. Indeed, suppose party $A$ has a fixed vote share $V>0.75$. Then from Theorem 3 (and Figure 4), as party $A$’s number of seats won decreases, the minimum value that the Mean-Median Difference can take on increases. In other words, as party $A$ wins fewer seats, the Mean-Median Difference is less able to show that party $A$ is winning an extremely small number of districts. In fact, once the seat share decreases to small enough values, the Mean-Median Difference is only able to take on values indicating a gerrymander for party $A$ (this can be seen in the “red crescent” in Figure 4) [^10] By symmetry, for a fixed vote share $V<0.25$, once the seat share increases large enough, the Mean-Median Difference is only able to take on values indicating a gerrymander against party $A$..
Similarly, for a fixed seat share $S$, the Mean-Median Difference does not allow for more extreme values indicating a less favorable position for party $A$ as party $A$’s vote share $V$ increases, nor does it allow for more extreme values indicating a more favorable position for party $A$ as $V$ decreases. For example, consider a fixed $S>0.5$. From Theorem 3 (and Figure 4), we see that as party $A$’s vote share $V$ increases, the Mean-Median Difference’s minimum value decreases, but only to a point. Then the Mean-Median Difference’s minimum value increases. That is, for large enough vote share $V$, if the seat share stays the same but the vote share increases, the Mean-Median Difference is less able to detect that party $A$ is winning extremely few districts. If the vote share increases large enough, the Mean-Median Difference can only indicate a gerrymander favoring party $A$ (as indicated by the red crescent in Figure 4) [^11] By symmetry, if $S<0.5$, for small enough vote share, if the seat share stays the same but the vote share decreases, the Mean-Median Difference is less able to detect that party $A$ is winning an extreme number of districts. And if the vote share decreases to be sufficiently small, the Mean-Median Difference can only indicate a gerrymander against party $A$ (as indicated by the blue crescent in Figure 4)..If we again fix $S>0.5$, as $V$ decreases, the Mean-Median Difference’s maximum value increases to a point, but then decreases. That is, even though party $A$ is winning the same number of seats, as the vote share decreases, the Mean-Median Difference is eventually less able to show that party $A$ is winning an extreme number of districts [^12] By symmetry, for $S<0.5$, as the vote share increases, the Mean-Median Difference is eventually less able to show that party $A$ is winning extremely few districts..
4.3.4 Mean-Median Difference: Elections with Values Counter to Public Expectations
In Section 2.2, we mentioned many instances in which the Mean-Median Difference has been used (or its usage has been promoted) in an effort to avoid maps that allow for one party to win an extreme number of districts. In this section, we give sample elections along with their Mean-Median Difference calculations. We believe these give convincing evidence that the Mean-Median Difference does not always act in the ways suggested by many of the testimonies, apps, reports, and policies mentioned in Section 2.2.
From Figure 4, similar to the observations we saw in Figure 3 in Section 4.3.2, we can see that the same coordinate position $(V,S)$ can have extremely different hues in the two images. This corresponds to the fact that, for fixed $V$ and $S$ values, the Mean-Median Difference can vary wildly. For example, consider the election data in Table 4
District 1 2 3 4 5 6 7 8 9 10
Vote share Election 1
1% 51% 51% 51% 91% 91% 91% 91% 91% 91%
Election 2
49% 59% 59% 59% 59% 59% 59% 99% 99% 99%
Table 4: Two different elections with $S=90\%$ and $V=70\%$. Election 1 has $\text{MM}=0.21$, while Election 2 has $\text{MM}=-0.11$.
Although both Elections in Table 4 have a vote share of 70% and seat share of 90%, the Mean-Median Difference considers Election 1 to be an extreme gerrymander in favor of the party in question, while Election 2 is considered to be an extreme gerrymander against the party in question. The different values here indicate that the Mean-Median Difference believes that 9 out of 10 seats is extremely large in the first election, but not nearly large enough in the second.
Additionally, note in Figure 4 that for $V>\frac{1}{2}$, MM has the same minimal value, regardless of $S$. For example, consider the elections in Table 5.
District 1 2 3 4 5 6 7 8 9 10 11
Vote share Election 1
47% 47% 47% 47% 47% 51% 74.8% 74.8% 74.8% 74.8% 74.8%
Election 2
47% 51% 51% 51% 51% 51% 71.6% 71.6% 71.6% 71.6% 71.6%
Table 5: Two different elections with $V=60\%$ and $\text{MM}=-0.09$. Election 1 has $S=\frac{6}{11}\approx 54.5\%$, and Election 2 has $S=\frac{10}{11}\approx 90.9\%$.
In each of the elections in Table 5, the vote share is 60% and the Mean-Median Difference is -0.09, indicating that each is a gerrymander against the party in question, even though the seat shares vary wildly (from 54.5% to 90.9%). The Mean-Median Difference is indicating that, for each of the elections in Table 5 (which have the party in question winning 6 and 10 seats out of 11 respectively), an extremely small number of districts is won. Indeed, since $\text{MM}=-0.09$ for each of those elections, the Partisan Bias is indicating that 6 and 10 seats won out of 11 are equally extremely small.
The last observation we make in this section is that for a fixed $V$ in the feasible region with $\frac{1}{2}\leq S\leq 1$, $\frac{1}{2}\leq V\leq 1$, the largest maximum values occur near the line $S=2V-\frac{1}{2}$. For example, consider the election data in Table 6:
District 1 2 3 4 5 6 7 8 9 10
Vote share Election 1
1% 1% 1% 1% 82.67% 82.67% 82.67% 82.67% 82.67% 82.67%
Election 2
1% 55.44% 55.44% 55.44% 55.44% 55.44% 55.44% 55.44% 55.44% 55.44%
Table 6: Two different elections with $V=50\%$. Election 1 has seat share 60% and $\text{MM}\approx 0.33$, while Election 2 has seat share 90% and $\text{MM}\approx 0.05$.
In both of these elections the vote share is 50%. In Election 1, the seat share is lower (6 seats out of 10) and the Mean-Median Difference is approximately 0.33, while in Election 2 the seat share is higher (9 seats out of 10) and the Mean-Median Difference is approximately 0.05. Hence, according to the Mean-Median Difference, these elections are both biased towards the party in question. However, since MM is larger for Election 1 (when 6 seats are won), 6 out of 10 seats won is considered a more extreme gerrymander for the party in question than in Election 2 (when 9 out of 10 seats are won).
4.4 The Special Case of Metric Value 0